Jane wants to make an acid. She has two possible compounds she could add to 250.0 g of water to
make this acid – EITHER 234.5 g of calcium oxide OR 678.9 g of tetraphosphorus decaoxide (she won’t use both, just one).
a. Which compound should she use? Explain your answer (with words and full sentences), and write any balanced chemical reactions that would help you justify your claim.

b.
What is the chemical formula of the acid, and what is the maximum mass of the acid she could
theoretically produce when she does this reaction?

c. What is the EXCESS reactant, and how many grams of this reactant are leftover when the reaction is
completed?
1
Expert's answer
2015-11-09T08:45:46-0500
a. Calcium oxide is a basic oxide. Soluble in water, releasing energy, forming calcium hydroxide:
СаО + Н2О → Са(ОН)2
Thus, this material is not suitable for producing acid.
P4O10 actively reacts with water to produce phosphoric acid:
P4O10 + 6 H2O → 4 H3PO4

b. Molar mass P4O10 is 283.9 g/mol
The 678.9 g P4O10 contains 678.9 / 283.9 = 2.39 mol P4O10
The molar mass of Н2О is 18.0 g / mol
The 250.0 g of Н2О contains 250.0 / 18.0 = 13.89 mol Н2О
In accordance with the reaction equation, 1 mol P4O10 reacted with 6 moles of water.
Then, to turn into acid 2.39 mol Р4O10 need 2.39 * 6 = 14.34 mol of water. Jane has 13.89 mol of water..
13.89 mol Н2О react to 13.89 / 6 = 2.315 mol P4O10.
From 2,315 mol P4O10 produced 2,315 * 4 = 9.26 mol of H3PO4
The molar mass of H3PO4 equal to 98.0 g / mol Then the mass of the resulting phosphoric acid is 9.26 * 98.0 = 907.5 g

c. Excess P4O10 is 2.39 - 2.315 = 0.075 mol.
Mass excess P4O10 is 0.075 * 283.9 = 21.3g